shift summary for owl shift 6/26/99

Theodoros Alexopoulos (theoalex@fnal.gov)
Sat, 26 Jun 1999 07:49:37 -0500 (CDT)

Shift Summary for Owl shift, 6/26/99

On Shift: JimH, TA
Sydney until 4:00am

Hours of beam delivered at all intensities: 6.7 hours
Hours of beam delivered at nominal intensity: 6.5 hours
Hours of E832 data (special runs not included): 6.5 hours
Hours of special runs (laser scans, muon runs...): 0.10 hours

Accounting of lost time: stopping and starting runs.

========================================================================
Runs ended this shift:
Run 13736 - 9.9M events - BEAM_832
Run 13737 - 0.9M events - BEAM_MU_4E12
Run 13738 - 5.5M events - BEAM_832

Tapes brought to FCC this shift:
QKC046,QKC047,QKC048,
QKS046,QKS047,QKS048,
QKK046,QKK047,QKK048,
QKB015,QKB014

========================================================================
Accesses during shift: a 10 mins access to check
the MU3X (channel 21 to 40) fluke power supply,
during MI power supply problem.

========================================================================
Overview of shift:

The shift was standard BEAM_832 running + a MUON run.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13736 00:10 run in progress

We investigated some more
the NM2SEED2 shift that was reported
in the evening shift summary.
Running data from the sdaq swic database,
we found that the problem
occured for runNum=13735, between spills
25 to 65. See Log XX, p.57 for more information.
02:38 run ended

13737 03:00 start muon run
03:10 end muon run, switch NM4AN to negative polarity
Sydney remade the L3 driver at RK request.
13738 03:30 Start BEAM_832. Sdaq gives MU3X power supply alarm;
also the illumination MU3X shows problems,
indicating a HV problem (see Log XX, p.60).
Paged Ed, he suggested
keep running, until we get a chance to get a
free controlled access to check the fluke power supply
of MU3X (for channels 21 to 40).
05:36 ktev2a alarm (97% full); remove some old files from
/ktev2a/CSI and /ktev2a/VETO
06:10 MI power supply problems, Est 15 mins for beam
06:40 Experts enroute for the MI power supply problem.
07:00 JimH and TA on control access to chek the
MU3X power supply. Found that the power supply
is at 2,500V although the front panel
switches are at 2,200V!!! Recycle the power
of this power supply, fixes the problem.
We should keep checking it...
07:15 stop 13738

========================================================================
Follow up Notes on Previous Shift

o Took the requsted muon run.

========================================================================
Items to Follow up on Next Shift

o If there is no beam for more than an hour, a laser scan
should be taken.